solve the quadratic equation by factorization
3x²-5x+2=0 ?
and can i be shown a checking method to show this is correct thank you

Answer by rothauserc(4718)

3x^2 -5x +2 = 0
:
(3x-2) * (x-1) = 0
*******************************
solutions are x = 2/3 and x = 1
*******************************
:
to check the answers, we substitute in the given equation
:
3x^2 -5x +2 = 0
:
1) x = 2/3
3(2/3)^2 -5(2/3) +2 = 0
4/3 -10/3 +2 = 0
-2 +2 = 0
0 = 0
:
2) x = 1
3(1)^2 -5(1) + 2 = 0
3 -5 +2 = 0
5 - 5 = 0
0 = 0
:
our answers check :-)
: